### v1

# InverseMatrixVT3D: An Efficient Projection Matrix-Based Approach for 3D Occupancy Prediction

Abstract
This paper introduces InverseMatrixVT3D, an efficient method for transforming multi-view image features into 3D feature volumes for 3D semantic occupancy prediction. Existing methods for constructing 3D volumes often rely on depth estimation, device-specific operators, or transformer queries, which hinders the widespread adoption of 3D occupancy models. In contrast, our approach leverages two projection matrices to store the static mapping relationships and matrix multiplications to efficiently generate global Bird’s Eye View (BEV) features and local 3D feature volumes. Specifically, we achieve this by performing matrix multiplications between multi-view image feature maps and two sparse projection matrices. We introduce a sparse matrix handling technique for the projection matrices to optimise GPU memory usage. Moreover, a global-local attention fusion module is proposed to integrate the global BEV features with the local 3D feature volumes to obtain the final 3D volume. We also employ a multi-scale supervision mechanism to further enhance performance. Comprehensive experiments on the nuScenes dataset demonstrate the simplicity and effectiveness of our method. The code will be made available at:https://github.com/DanielMing123/InverseMatrixVT3D
Figure 1: The figure above shows the pipeline of three approaches: Depth-Estimation-based approach (upper), Query-based approach (middle), and our approach (lower). We simplify the generation of the 3D Feature Volume by adopting a matrix multiplication method.
I Introduction
Understanding the surrounding scene’s three-dimensional (3D) geometry is fundamental to developing autonomous driving (AV) systems. While lidar-based methods [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] which utilize explicit depth measurements have been performing exceptionally well on public datasets [14, 15, 16, 17], they are hindered by the expensive cost of sensors and the sparsity of data points. As a result, the broader application of lidar-based methods is limited.
In recent years, vision-centric AV systems have garnered significant attention as a promising strategy due to its cost-effectiveness, stability, and generality. By utilizing multi-camera images as inputs, this approach has demonstrated competitive performance across various 3D perception tasks, including depth estimation [18, 19, 20], 3D object detection [21, 22, 23, 24, 25, 26, 27, 28], online high-definition (HD) map construction [29, 30, 31, 32, 33], and semantic map construction [34, 35, 36].
3D object detection based on fusing surround view cameras can play a crucial role in 3D perception. However, it faces challenges in handling new scenarios. One of the challenges is the finite number of semantic classes in the training dataset, making it difficult to create a model for every potential scenario that might be encountered on the road. In contrast, a more practical approach to depict the vehicle’s surrounding environment is by directly reconstructing the 3D scenes. In pursuit of this objective, several methods [37, 38, 39, 40, 41] have been investigated to predict the 3D occupancy of a scene directly. These methods involve voxelizing the 3D space and assigning a probability to each voxel to determine its occupancy state–whether it is occupied or not. We argue that 3D occupancy serves as an adequate representation of the vehicle’s surrounding environment. This representation inherently ensures geometric consistency and can accurately describe occluded areas. Furthermore, it is more robust towards object classes that do not exist in the training dataset. Despite the promise of these methods, their inner structure is quite complex, and some methods require additional sensors to provide supervision signals. For instance, the method proposed in [40] relies on a lidar sensor to enhance performance through depth estimation supervision (Figure 1, upper). Additionally, the methods proposed in [38, 39] extensively employ query-based modules [42, 43, 44] to aggregate image features for the final 3D feature volume (Figure 1, middle).
To efficiently and effectively represent a 3D scene using 3D occupancy, we propose InverseMatrixVT3D (Figure 1, bottom). Our method focuses on constructing projection matrices and simplifying the generation of local 3D feature volumes and global Bird’s Eye View (BEV) features through matrix multiplication between multi-scale feature maps and projection matrices. Additionally, we employ a sparse matrix handling technique to optimize GPU memory usage when using these sparse projection matrices. Furthermore, we introduce a global-local fusion module to integrate global BEV features with local 3D feature volumes, resulting in the final 3D volume. We also apply a multi-scale supervision mechanism to each level to further enhance performance. Through comparisons with state-of-the-art (SOTA) algorithms on the nuScenes benchmark, we demonstrate our method’s simplicity and superior overall performance.
The main contributions of this paper are summarized below.
  • The novel projection matrix-based approach is proposed to construct local 3D feature volumes and global BEV features.
  • A global-local fusion module is proposed to integrate the information, resulting in the final 3D volume.
  • We compare our approach with other state-of-the-art (SOTA) algorithms in the 3D semantic occupancy prediction task to prove the superior performance of our method.
The remainder of this paper is structured as follows: Section II provides an overview of related research and identifies the key differences between this study and previous publications. Section III outlines the general framework of InverseMatrixVT3D and offers a detailed explanation of the implementation of each module. Section IV presents the findings of our experiments. Finally, Section V provides the conclusion of our work.
II Related Work
II-A Depth-Estimation-Based BEV Perception
Lift-Splat-Shoot (LSS) [21] introduced a lift operation for depth estimation using multi-view cameras. This process generates a pseudo-3D point cloud by merging the pseudo-3D points obtained from the multi-view images. The pseudo-3D point cloud is projected onto the ego vehicle coordinate system by leveraging intrinsic and extrinsic parameters. Subsequently, a splat operation is employed to accumulate features in the same BEV plane grid with varying z-direction heights, resulting in the final BEV features. This study pioneered the generation of BEV features through depth estimation and the fusion of multi-camera information in BEV perception. Another important contribution in this area was made by BEVDepth [24], highlighting the crucial impact of depth estimation precision on the model’s performance. To address this, BEVDepth introduced a real lidar-sensor based 3D point cloud as an additional supervision signal for the depth estimation stage. Furthermore, BEVStereo [25] improved depth estimation performance by introducing the temporal stereo method.
Figure 2: Overall architecture of InverseMatrixVT3D. Firstly, the multi-camera images were inputted into the 2D backbone network to extract features at multiple scales. Subsequently, a multi-scale global local projection module was employed to construct multi-scale 3D feature volumes and BEV planes. A global-local attention fusion module was applied to each 3D feature volume and BEV plane at every level to obtain the final 3D feature volume. Finally, the 3D volume at each level was upsampled using 3D deconvolution for skip-connection, and a supervision signal was also applied at each level.
II-B Depth-Estimation Based 3D Semantic Occupancy Prediction
Based on the success of depth-estimation based BEV perception algorithms, several works [45, 46, 41, 40] have focused on constructing the 3D feature volume using pseudo-3D points. These approaches replace the previous splat operation, which generates the BEV feature, with a voxel-pooling operation. This new approach voxelizes the pseudo-3D point cloud and proposes several refinement modules to enhance the 3D feature volume. OCCFormer [40] introduces a dual-path transformer block to refine the BEV slice of the 3D feature volume, enhancing the long-range modelling capability of their model. FB-OCC [41] proposes an additional backward view-transformation module to improve the semantic information in the final 3D feature volume. Multi-Scale Occ [46] leverages a multi-scale fusion mechanism to capture global and local detail information in the 3D feature volume.
While depth-estimation-based approaches have achieved remarkable performance, they have one significant drawback: requiring depth ground truth labels to boost depth-estimation performance, which in turn boosts the model’s overall performance. This requirement introduces extra effort during the training process. In this paper, we propose a method that eliminates the need for depth estimation by solely relying on multi-view images to construct a 3D feature volume. Our approach achieves superior performance compared to depth-estimation based approaches.
II-C Query-Based BEV Perception
The BEVFormer series [26, 47] is one of the pioneering query-based approaches to BEV perception. By defining a BEV query plane and utilizing intrinsic and extrinsic parameters, the query vector in each grid of the BEV plane can be projected onto multi-view images, enabling the aggregation of features and the generation of the final BEV feature. The temporal self-attention module also leverages historical BEV features to enhance the model’s performance further. Similarly, the DETR3D [48] predefines a set of 3D query points as reference points. The intrinsic and extrinsic parameters are then used to project those points onto multi-view images to aggregate features. These works successfully eliminate the need for depth estimation by employing transformer blocks and relying solely on multi-view image features for BEV perception.
II-D Query-Based 3D Semantic Occupancy Prediction
Building upon the success of query-based BEV perception algorithms, TPVFormer [38] introduces an extension to the BEV query by encompassing three perpendicular planes. This approach aims to capture the 3D world from multiple orthogonal perspectives. Similarly, SurroundOcc [39] further expands the three-perpendicular plane concept to a 3D query volume. Leveraging intrinsic and extrinsic parameters, each query vector is projected onto multi-view images to aggregate dense features. Additionally, PanoOcc [49] proposes an efficient method for processing the dense 3D feature volume using a sparse representation approach.
Given the success of query-based approaches, it is important to highlight that the extensive use of transformer blocks in these methods often results in slow and inefficient training processes and high GPU memory consumption. In contrast, our proposed method eliminates the requirement for transformer-based querying and depth estimation while constructing the 3D feature volume. Our approach significantly improves model efficiency and enhances overall performance.
III InverseMatrixVT3D
In this paper, our objective is to generate a dense 3D occupancy grid of the surrounding scene using multi-camera images $Img=\left\{Img^{1},Img^{2},\cdots,Img^{N}\right\}$. Formally, the problem can be described as follows:
$3DOcc=VT(Img^{1},Img^{2},\dots,Img^{N})$ (1)
where VT is the neural network that leverages the projection matrix to aggregate features for 3D occupancy, the final 3D occupancy prediction result denoted as $3DOcc\in R^{X\times Y\times Z}$, represents the semantic property of the grids and has values ranging from 0 to 16. In our case, a class value of 0 indicates that the grid is empty.
Figure 2 exhibits the overall architecture of our method. Initially, given a set of surrounding multi-camera images, we first use a 2D backbone network(e.g.ResNet101-DCN) to extract $Nc$ cameras and $L$ levels multi-scale features $X=\left\{\left\{X_{n}^{l}\right\}_{n=1}^{Nc}\in R^{C_{l}\times H_{l}\times W_{l}}\right\}_{l=1}^{L}$. For each level, we construct two projection matrices, namely, global projection matrix $VT\_XY^{l}\in R^{(Nc\times H_{l}\times W_{l})\times(X_{l}\times Y_{l})}$ and local projection matrix $VT\_XYZ^{l}\in R^{(Nc\times H_{l}\times W_{l})\times(X_{l}\times Y_{l}\times Z_{l})}$. The feature maps at each level are multiplied with these projection matrices, resulting in the 3D feature volume $F_{local}^{l}\in{R}^{C_{l}\times X_{l}\times Y_{l}\times Z_{l}}$ and the Bird’s-eye view (BEV) feature $F_{global}^{l}\in{R}^{C_{l}\times X_{l}\times Y_{l}}$. Subsequently, the global-local attention fusion module merges information from these two features, producing the final 3D volume. Furthermore, the 3D volume at each level is upsampled using 3D deconvolution and integrated with the higher-level 3D volume through skip-connection. Finally, the dense occupancy ground truth from [39] is applied to supervise the 3D volume at each level with a decayed loss weight.
III-A Multi-Camera Images Feature Extractor
The purpose of the multi-camera image feature extractor is to capture both spatial and semantic features from the surrounding perspective view. These extracted features are the basis for the subsequent 3D occupancy prediction task. In our approach, we first employ a 2D backbone network to extract multi-scale feature maps. Then, a feature-pyramid network (FPN) is followed to further fuse feature output from different stages of the 2D backbone. The resulting feature maps have resolutions that are $\frac{1}{8}$, $\frac{1}{16}$, and $\frac{1}{32}$ of the input image resolution, respectively. The feature maps with smaller resolutions contain abundant semantic information, which assists the model in predicting the semantic class of each voxel grid. Conversely, the feature maps with larger resolutions provide richer spatial details and better guide the model regarding whether the current voxel grid is occupied or unoccupied.
III-B Global and Local Projection Matrix Generation
Figure 3: The predefined 3D volume feature sampling process. Each voxel grid is initially divided into $N^{3}$ subspaces along the vertical and horizontal directions, and the centre of each subspace serves as a sample point. These sample points are then projected onto all the multi-view feature maps to aggregate the corresponding features for the voxel grid.
In our approach, constructing global and local projection matrices is critical in gathering information for the local 3D feature volume and the global BEV feature. This differs from the method used in Occformer[40], which relies on depth estimation and voxel pooling, or the methods employed in TPVFormer[38] and SurroundOcc[39], which involve transformer-based query. To begin with, we establish a set of predefined 3D volume spaces denoted by $F_{3D}^{l}\in{R}^{X^{l}\times Y^{l}\times Z^{l}}$ for each level of multi-view feature maps under the ego vehicle’s coordinate system. The ego vehicle is positioned at the center of these 3D volumes. Within each 3D volume, we divide the voxel grid equally into $N^{3}$ subspaces along the horizontal and vertical directions. Each center of the subspace serves as a sample point. Thus, for each level of the 3D volume space, we have a total of $X^{l}\times Y^{l}\times Z^{l}\times N^{3}$ sample points. Next, we project each sample point from each voxel grid onto the corresponding level of multi-view feature map $X^{l}$ using extrinsic and intrinsic parameters. Sample points that fall outside the boundaries of the feature maps or generate negative depth values are filtered out. The features hit by each sample point are then aggregated, resulting in a feature vector representing the corresponding voxel grid (see Figure 3).
The feature sampling process is a static mapping and can be represented by constructing the projection matrices $VT\_XYZ^{l}\in R^{(Nc\times H_{l}\times W_{l})\times(X_{l}\times Y_{l}\times Z_{l})}$ as exhibited in Figure 4. Moreover, height information can be compressed further by constructing the projection matrix $VT\_XY^{l}\in R^{(Nc\times H_{l}\times W_{l})\times(X_{l}\times Y_{l})}$ as shown in Figure 5. As a result, the generation of both the local 3D feature volume and global BEV feature can be greatly simplified as a matrix multiplication between the multi-view feature maps and the two projection matrices:
$F_{local}^{l}=X^{l}\cdot VT\_XYZ^{l}$ (2)
and
$F_{global}^{l}=X^{l}\cdot VT\_XY^{l}$ (3)
where $X^{l}\in R^{N_{c}\times H_{l}\times W_{l}}$ represents the multi-view feature maps at the $l$-th level.
During the construction process of the global and local projection matrices, we observed that these matrices exhibit extensive sparsity. Consequently, the GPU memory utilization for constructing these matrices increases exponentially with their resolution. To optimize GPU memory utilization, we utilize the compressed sparse row (CSR) technique [50]. This technique stores only the non-zero values and their associated indices when constructing and storing the sparse matrices. By applying this technique, we can dramatically decrease the GPU memory usage for our highest 3D volume resolution from 15GB to 200MB.
Figure 4: The local projection matrix VT_XYZ represents the static mapping relationship for the feature sampling process of the local 3D feature volume.
Figure 5: The global projection matrix VT_XY represents the static mapping relationship for the feature sampling process of the global Bird’s Eye View (BEV) feature.
III-C Global Local Attention Fusion
To enhance the ability of the final 3D feature volume to capture both global and local details, we introduce the Global-Local Attention Fusion module. The detailed structure of the Global-Local Attention Fusion module is depicted in Figure 6.
Figure 6: The global-local fusion module. The global BEV feature and the local 3D feature volume are refined using traditional convolutional layers. The global BEV feature undergoes additional enhancement through an efficient window attention module and a bottleneck ASPP module. It then merges with the local 3D feature volume, resulting in the final 3D volume.
Inspired by [51], the module begins by applying traditional 2D and 3D convolution operations to enhance the global BEV feature and the local 3D feature volume. Building on recent advancements that emphasize the significance of locality and efficiency in transformers [52, 53] and the importance of the BEV plane [40], we incorporate an efficient window attention module from [52] and a bottleneck ASPP module to further refine the global BEV feature, enabling it to capture long-range global semantic information. Simultaneously, the local 3D feature volume is processed through the Feed-Forward Network (FFN) to generate attention weights. The BEV feature, refined by the window attention and bottleneck ASPP module, undergoes dimension expansion and is element-wise multiplied by the attention weights. This resulting feature is then added to the refined local 3D feature volume to produce the final 3D feature volume. The whole process can be described as follows:
$F_{3DOcc}^{l}=F_{local}^{l}+\sigma(FNN(F_{local}^{l}))\cdot Expand(F_{global}^{l})$ (4)
where $\sigma()$ refers to applying the sigmoid function to the output of the FFN; this function constrains the attention weight to the [0,1] range. Additionally, the Expand operation refers to expanding the dimension of $F_{global}^{l}$ to match the dimension of $F_{local}^{l}$.
IV Experimental Results
IV-A Implementation Details
The InverseMatrixVT3D uses ResNet101-DCN [54, 55] as a 2D backbone with a checkpoint from FCOS3D [56] to extract image features. The features of stages 1,2,3 of the backbone are fed to FPN [57], resulting in 3-level multi-scale image features. The network architecture comprises four levels ($L=4$), with no skip connection applied to the highest level. For the paths corresponding to levels 1, 2, and 3, we employ divided schemas of N=3, 4, and 5, respectively, to create sets of sample points. The AdamW optimizer with an initial learning rate of 5e-5 and weight decay of 0.01 is employed for optimization. The learning rate is decayed using a multi-step scheduler. For the 3D semantic occupancy prediction task, following the approach presented in [39, 37], we incorporate focal loss [58], Lovasz-softmax loss [59] and scene-class affinity loss [37] to handle the significant sparsity of free space in the ground truth dataset. Regarding data augmentation, random resize, rotation and flip operations are applied in the image space, following established practices for BEV-based 3D object detection [23, 24, 25, 26] and the compared methods [38, 39, 40, 37]. Our model includes two resolution settings. The predicted occupancy for evaluation is specified at a resolution of $256\times 256\times 32$ and $200\times 200\times 16$, respectively. We further fixed the projection matrices under two resolution settings during testing to make the comparison. The model is trained using eight A10G GPUs with 24GB of memory, and it has been trained for 24 epochs on 3D semantic occupancy prediction tasks.
IV-B Dataset
The nuScenes dataset [14], a vast autonomous driving dataset, serves as the data source for our experiments. As the test set lacks semantic labels, we train our model on the training set and assess its performance using the validation set. We set the range for occupancy prediction as [-50, 50] meters for the X and Y axes and [-5, 3] meters for the Z axis. The input images have a $1600\times 900$ pixels resolution, while the final output occupancy is represented with a resolution of $256\times 256\times 32$ for the base version and $200\times 200\times 16$ for the small base version. We have conducted experiments on 3D semantic occupancy prediction tasks to provide quantitative results. The dense labels used for the 3D semantic occupancy prediction task are sourced from [39]. Additionally, we provide qualitative visualizations of the results for the 3D semantic occupancy prediction task.
IV-C Performance Evaluate Metrics
To assess the performance of various state-of-the-art (SOTA) algorithms and compare them with our approach in the 3D semantic occupancy prediction task, we utilize the intersection over union (IoU) to evaluate each semantic class. Moreover, we employ the mean IoU overall semantic classes (mIoU) as a comprehensive evaluation metric:
$IoU=\frac{TP}{TP+FP+FN}$ (5)
and
$mIoU=\frac{1}{Cls}\sum_{i=1}^{Cls}\frac{TP_{i}}{TP_{i}+FP_{i}+FN_{i}}$ (6)
where $TP$, $FP$, and $FN$ represent the counts of true positives, false positives, and false negatives in our predictions, respectively, while $Cls$ denotes the total class number.
IV-D Model Performance Analysis
We perform the task of multi-camera 3D semantic occupancy prediction on the nuScenes dataset. To evaluate the performance of our proposed InverseMatrixVT3D, we compare it with SOTA algorithms and present the results in Table I. All methods are trained using dense 3D occupancy ground truth labels obtained from [39]. Our model has two different settings for two resolutions: the InverseMatrixVT3D(Base) and InverseMatrixVT3D(Small Base), which includes the projection matrices generation process, and the InverseMatrixVT3D(Fix) and InverseMatrixVT3D(Small Fix), which fixes the projection matrices. Our model displays remarkable performance in both settings in predicting crucial dynamic objects on the road, including buses, cars, construction vehicles, and trucks. Additionally, our model excels in identifying diverse background objects, such as drivable surfaces, terrain, and vegetation. These results unequivocally highlight our model’s superior 3D world modelling capability compared to other SOTA algorithms.
Our model aggregates features for the final 3D volume based on sampling at specific sampling locations. Compared to the transformer-based approach, whose sampling locations can be adjusted based on the query vector, our approach’s sampling locations are fixed. Thus, when dealing with some small objects such as traffic cones and bicycles, our method aggregates fewer features from the image and is inferior to SurroundOcc in predicting those classes. However, our sampling density is significantly larger than the transformer-based approach, enabling our method to make good predictions on relatively large objects in the scene. In summary, our method sacrifices sampling flexibility to improve the sampling density. This underlying principle adjustment boosts model performance significantly in predicting relatively large objects in the scene.
Method Resolution 3DOcc mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation MonoScene [37] $200\times 200\times 16$ 7.31 4.03 0.35 8.00 8.04 2.90 0.28 1.16 0.67 4.01 4.35 27.72 5.20 15.13 11.29 9.03 14.86 Atlas [60] $200\times 200\times 16$ 15.00 10.64 5.68 19.66 24.94 8.90 8.84 6.47 3.28 10.42 16.21 34.86 15.46 21.89 20.95 11.21 20.54 BEVFormer [26] $200\times 200$ 16.75 14.22 6.58 23.46 28.28 8.66 10.77 6.64 4.05 11.20 17.78 37.28 18.00 22.88 22.17 13.80 22.21 TPVFormer [38] $200\times 200\times 16$ 17.10 15.96 5.31 23.86 27.32 9.79 8.74 7.09 5.20 10.97 19.22 38.87 21.25 24.26 23.15 11.73 20.81 SurroundOcc [39] $200\times 200\times 16$ 20.30 20.59 11.68 28.06 30.86 10.70 15.14 14.09 12.06 14.38 22.26 37.29 23.70 24.49 22.77 14.89 21.86 InverseMatrixVT3D (Base) $256\times 256\times 32$ 26.03 26.15 9.76 42.27 41.89 17.83 14.74 22.88 10.30 21.78 36.90 50.82 21.15 24.45 26.76 16.86 32.03 InverseMatrixVT3D (Small Base) $200\times 200\times 16$ 25.95 26.13 15.19 42.65 43.92 11.42 20.82 22.68 13.45 22.33 36.04 47.47 16.26 21.47 25.33 16.37 33.73 InverseMatrixVT3D (Fix) $256\times 256\times 32$ 24.58 23.19 7.86 43.17 43.16 16.80 16.03 18.16 8.81 24.67 35.93 41.60 17.68 21.11 21.54 22.06 31.45 InverseMatrixVT3D (Small Fix) $200\times 200\times 16$ 22.68 20.91 12.06 38.52 35.53 10.53 21.49 20.12 12.36 20.49 30.31 36.23 14.60 19.25 22.78 15.89 31.76
TABLE I: 3D semantic occupancy prediction results on nuScenes validation set. All methods are trained with dense occupancy labels from [39].
IV-E Challenging Scenes Qualitative Analysis
Figure 6a presents a scene where multiple objects of different semantic classes overlap and crowd. Despite this complexity, our method accurately predicts each object’s semantic class and 3D spatial location. Furthermore, the edges between objects of different classes are delineated. Figure 6b exhibits the scene where a construction vehicle is severely obscured by a large truck while multiple traffic cones and barriers crowd and overlap. Even under such a complex scenario, our model can still infer the general occupancy status of the construction vehicle and generate a meticulous 3D occupancy prediction with distinct boundaries.
We further demonstrate our model’s powerful 3D modelling capability by presenting the prediction results in challenging rainy and night-time scenes, as shown in Figure 8. In each subfigure, the prediction result generated by our model is displayed on the bottom left, while the ground truth is shown on the bottom right. Regarding the rainy scenario depicted in Figure 7a, we observed an inconsistency in the ground truth labels. This discrepancy is attributed to the lidar’s poor performance during rainy weather due to the reflection effect. Consequently, the ground truth labels become inconsistent. However, our model’s prediction result is remarkably accurate. It not only fills in the missing information from the ground truth but also successfully detects a pedestrian crossing the road (indicated by the red circle), even though the person is located at a significant distance from the ego vehicle. Moreover, in the night-time scene illustrated in Figure 7b, our model effectively predicts the presence of a motorcycle (marked by the red circle) in the distance ahead despite the low ambient light conditions.
We also present scenes showcasing crowded multiple objects and rainy day crossroads in Figure 9. The left column displays the multi-camera images, the middle column shows the prediction results of our model, and the right column exhibits the ground truth labels. In the scenarios with crowded multiple objects represented in the first and second rows of Figure 9, our model accurately segments each object. Furthermore, in the rainy city intersection scenario depicted in the last row of Figure 9, our model successfully predicts the road structure even in the presence of missing information in the ground truth labels.
a Highly overlapping scene
b Severely obscured scene
$\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ construction vehicle $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck
$\bullet$ driveable surface $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation $\bullet$ ego vehicle
Figure 7: Challenging scenes qualitative analysis. Despite the presence of multiple objects that overlap or severely obscure each other, our method is capable of accurately predicting detailed occupancy and delineating their edges. Better viewed when zoomed in.
a Rainy scene
b Night-time scene
$\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ construction vehicle $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck
$\bullet$ driveable surface $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation $\bullet$ ego vehicle
Figure 8: Rainy and night-time scene. Despite challenging lighting conditions, our approach successfully predicts moving objects with high accuracy, even when they are far away from the ego vehicle. Better viewed when zoomed in.
a Multi-Camera Images
b Prediction Result
c Ground-Truth
$\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ construction vehicle $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck
$\bullet$ driveable surface $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation $\bullet$ ego vehicle
Figure 9: Challenge scenes. Despite multiple different objects being crowded or missing information in the ground truth, our model can still accurately segment each object.
IV-F Model Efficiency
Method Latency (s) ($\downarrow$) Memory (GB) ($\downarrow$) NeWCRFs [61] 1.07 14.5 MonoScene [37] 0.87 20.3 Adabins [62] 0.75 15.5 SurroundDepth [18] 0.73 12.4 SurroundOcc [39] 0.34 5.9 TPVFormer [38] 0.32 5.1 BEVFormer [26] 0.31 4.5 InverseMatrixVT3D (Base) 0.68 10.2 InverseMatrixVT3D (Small Base) 0.5 7.8 InverseMatrixVT3D (Fix) 0.45 4.88 InverseMatrixVT3D (Small Fix) 0.34 4.82
TABLE II: Model efficiency comparison of different methods. The experiments are performed on a single RTX 3090 using six multi-camera images. For input image resolution, all methods adopt $1600\times 900$. $\downarrow$:the lower the better.
Table II presents a comparison of the inference time and inference memory among different methods. The experiments are conducted on a single RTX 3090 using six multi-camera images. All methods adopt an image resolution of $1600\times 900$. The two base version models, which include the generation of projection matrices, have runtimes of 0.68 and 0.5 seconds for a single data sample. However, by fixing the projection matrices and removing the generation procedure, the fixed version model achieves exceptional real-time performance due to the utilization of projection matrices that facilitate 3D volume generation.
IV-G Ablation Study
IV-G1 Global Local Attention Fusion
We perform an ablation study on the global-local attention fusion module, and the experiment results are presented in Table III. When we remove the fusion module and directly use the local 3D feature volume as the final 3D volume, we observe around 1.9% degradation in the model’s performance. This degradation could be attributed to the inability of the final 3D volume to capture long-range global semantic information.
Method 3DOcc mIoU w fusion module 23.19 w/o fusion module 21.32
TABLE III: Ablation study for global-local attention fusion module
IV-G2 Multi-Scale Supervision
We conducted an ablation study on the multi-scale supervision mechanism, and the experiment results are presented in Table IV. The upsample and skip connection structure remains in the model where the multi-scale supervision mechanism is removed, but we only utilize the highest-level 3D feature volume for supervision. The experimental results demonstrate that the coarse-to-fine supervision schema can enhance performance by approximately 2.2%.
Method 3DOcc mIoU w Multi-Scale Supervision 23.19 w/o Multi-Scale Supervision 21.07
TABLE IV: Ablation study for Multi-Scale Supervision
IV-G3 Dividing Schemas For Each Level
We performed an ablation study on the division schemas for each level, and the experiment results are presented in Table V. We used different N settings for levels 1, 2, and 3 to generate sample points, and the experimental results aligned with our expectations. Specifically, when the divide setting N is larger, more sample points can be generated, leading to dense sampling and rich feature aggregation. We also observed a trend where a larger divide setting at higher levels has a more significant impact on the final model performance, indicating the increased importance of higher-level feature aggregation.
N setting of level 1,2,3 3DOcc mIoU $[3,4,5]$ 23.19 $[2,3,4]$ 20.01 $[1,2,3]$ 17.24 $[1,4,6]$ 19.33
TABLE V: Ablation study of dividing schemas for each level
V Conclusion
In this paper, we propose InverseMatrixVT3D, a method for vision-centric 3D semantic occupancy prediction. Our approach leverages predefined sample points for each scale of 3D volumes and constructs projection matrices to represent the fixed sampling process. Through matrix multiplication between multi-view feature maps and projection matrices in a multi-scale fashion, we generate local 3D feature volumes and global BEV features. These features are further merged using our proposed global-local fusion module, resulting in the final 3D volume at each level. Lastly, the 3D volumes at each level are upsampled and fused using a 3D deconvolution layer. Unlike other SOTA algorithms, our approach does not require depth estimation or transformer-based query, making the 3D volume generation process simple and efficient. Experimental results on nuScenes datasets demonstrate the superiority of our method in benchmark comparisons.

### v2

@@
 Specifically, we achieve this by performing matrix multiplications between multi-view image feature maps and two sparse projection matrices.
-We introduce a sparse matrix handling technique for the projection matrices to optimise GPU memory usage.
+We introduce a sparse matrix handling technique for the projection matrices to optimize GPU memory usage.
 Moreover, a global-local attention fusion module is proposed to integrate the global BEV features with the local 3D feature volumes to obtain the final 3D volume.
-We also employ a multi-scale supervision mechanism to further enhance performance.
-Comprehensive experiments on the nuScenes dataset demonstrate the simplicity and effectiveness of our method.
-The code will be made available at:https://github.com/DanielMing123/InverseMatrixVT3D
-Figure 1: The figure above shows the pipeline of three approaches: Depth-Estimation-based approach (upper), Query-based approach (middle), and our approach (lower).
-We simplify the generation of the 3D Feature Volume by adopting a matrix multiplication method.
+We also employ a multi-scale supervision mechanism to enhance performance further.
+Extensive experiments performed on the nuScenes and SemanticKITTI datasets reveal that our approach not only stands out for its simplicity and effectiveness but also achieves the top performance in detecting vulnerable road users (VRU), crucial for autonomous driving and road safety.
+The code has been made available at: https://github.com/DanielMing123/InverseMatrixVT3D
 I INTRODUCTION
 Understanding the surrounding scene’s three-dimensional (3D) geometry is fundamental to developing autonomous driving (AV) systems.
-While lidar-based methods [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] which utilize explicit depth measurements have been performing exceptionally well on public datasets [14, 15, 16, 17], they are hindered by the expensive cost of sensors and the sparsity of data points.
+While lidar-based methods that utilize explicit depth measurements have been performing exceptionally well on public datasets [1, 2], they are hindered by the expensive cost of sensors and the sparsity of data points.
 As a result, the broader application of lidar-based methods is limited.
-In recent years, vision-centric AV systems have garnered significant attention as a promising strategy due to its cost-effectiveness, stability, and generality.
-By utilizing multi-camera images as inputs, this approach has demonstrated competitive performance across various 3D perception tasks, including depth estimation [18, 19, 20], 3D object detection [21, 22, 23, 24, 25, 26, 27, 28], online high-definition (HD) map construction [29, 30, 31, 32, 33], and semantic map construction [34, 35, 36].
-3D object detection based on fusing surround view cameras can play a crucial role in 3D perception.
+Vision-centric AV systems have garnered significant attention in recent years as a promising strategy due to its cost-effectiveness, stability, and generality.
+By utilizing multi-camera images as inputs, this approach has demonstrated competitive performance across various 3D perception tasks, including depth estimation [3, 4], 3D object detection [5, 6, 7], online high-definition (HD) map construction [8, 9, 10], and semantic map construction [11, 12].
+Figure 1: Pipeline of three approaches: Depth-Estimation-based approach (upper), Query-based approach (middle), and our approach (lower).
+We simplify the generation of the 3D Feature Volume by adopting a matrix multiplication method.
+3D object detection based on fusing surround view cameras can be crucial in 3D perception.
 However, it faces challenges in handling new scenarios.
@@
 In contrast, a more practical approach to depict the vehicle’s surrounding environment is by directly reconstructing the 3D scenes.
-In pursuit of this objective, several methods [37, 38, 39, 40, 41] have been investigated to predict the 3D occupancy of a scene directly.
+In pursuit of this objective, several methods [13, 14, 15] have been investigated to predict the 3D occupancy of a scene directly.
 These methods involve voxelizing the 3D space and assigning a probability to each voxel to determine its occupancy state–whether it is occupied or not.
@@
 Despite the promise of these methods, their inner structure is quite complex, and some methods require additional sensors to provide supervision signals.
-For instance, the method proposed in [40] relies on a lidar sensor to enhance performance through depth estimation supervision (Figure 1, upper).
-Additionally, the methods proposed in [38, 39] extensively employ query-based modules [42, 43, 44] to aggregate image features for the final 3D feature volume (Figure 1, middle).
-To efficiently and effectively represent a 3D scene using 3D occupancy, we propose InverseMatrixVT3D (Figure 1, bottom).
+For instance, the method proposed in [15] relies on a lidar sensor to enhance performance through depth estimation supervision (Fig.
+1, upper).
+Additionally, the methods proposed in [13, 14] extensively employ query-based modules [16] to aggregate image features for the final 3D feature volume (Fig.
+1, middle).
+To efficiently and effectively represent a 3D scene using 3D occupancy, we propose InverseMatrixVT3D (Fig.
+1, bottom).
 Our method focuses on constructing projection matrices and simplifying the generation of local 3D feature volumes and global Bird’s Eye View (BEV) features through matrix multiplication between multi-scale feature maps and projection matrices.
@@
 We also apply a multi-scale supervision mechanism to each level to further enhance performance.
-Through comparisons with state-of-the-art (SOTA) algorithms on the nuScenes benchmark, we demonstrate our method’s simplicity and superior overall performance.
+Through comparisons with other state-of-the-art (SOTA) algorithms on the nuScenes and SemanticKITTI benchmarks, we demonstrate our method not only excels in its simplicity and effectiveness but also achieves the best performance in detecting vulnerable road users (VRU), i.e.
+pedestrians, motorcycles, and bicycles, which is a critical task for autonomous driving and road safety.
 The main contributions of this paper are summarized below.
-  • The novel projection matrix-based approach is proposed to construct local 3D feature volumes and global BEV features.
-  • A global-local fusion module is proposed to integrate the information, resulting in the final 3D volume.
-  • We compare our approach with other state-of-the-art (SOTA) algorithms in the 3D semantic occupancy prediction task to prove the superior performance of our method.
+  • A novel projection matrix-based approach is proposed to simplify the local 3D feature volume and global BEV feature construction.
+  • A global-local fusion module is proposed to integrate global long-range information from the BEV feature with local spatial detail information from the 3D feature volume, resulting in the final 3D volume.
+  • We compare our approach with other state-of-the-art (SOTA) algorithms in the 3D semantic occupancy prediction task to prove the simplicity and effectiveness of our method.
 The remainder of this paper is structured as follows: Section II provides an overview of related research and identifies the key differences between this study and previous publications.
@@
 II Related Work
-II-A Depth-Estimation-Based BEV Perception
-Lift-Splat-Shoot (LSS) [21] introduced a lift operation for depth estimation using multi-view cameras.
-This process generates a pseudo-3D point cloud by merging the pseudo-3D points obtained from the multi-view images.
-The pseudo-3D point cloud is projected onto the ego vehicle coordinate system by leveraging intrinsic and extrinsic parameters.
-Subsequently, a splat operation is employed to accumulate features in the same BEV plane grid with varying z-direction heights, resulting in the final BEV features.
-This study pioneered the generation of BEV features through depth estimation and the fusion of multi-camera information in BEV perception.
-Another important contribution in this area was made by BEVDepth [24], highlighting the crucial impact of depth estimation precision on the model’s performance.
-To address this, BEVDepth introduced a real lidar-sensor based 3D point cloud as an additional supervision signal for the depth estimation stage.
-Furthermore, BEVStereo [25] improved depth estimation performance by introducing the temporal stereo method.
 Figure 2: Overall architecture of InverseMatrixVT3D.
@@
 Finally, the 3D volume at each level was upsampled using 3D deconvolution for skip-connection, and a supervision signal was also applied at each level.
-II-B Depth-Estimation Based 3D Semantic Occupancy Prediction
-Based on the success of depth-estimation based BEV perception algorithms, several works [45, 46, 41, 40] have focused on constructing the 3D feature volume using pseudo-3D points.
-These approaches replace the previous splat operation, which generates the BEV feature, with a voxel-pooling operation.
+II-A Depth-Estimation Based 3D Semantic Occupancy Prediction
+Based on the success of depth-estimation-based BEV perception algorithms, several works [17, 18, 19, 15] have focused on constructing the 3D feature volume using pseudo-3D points.
+These approaches replace the previous splat operation in LSS [20], which generates the BEV feature, with a voxel-pooling operation.
 This new approach voxelizes the pseudo-3D point cloud and proposes several refinement modules to enhance the 3D feature volume.
-OCCFormer [40] introduces a dual-path transformer block to refine the BEV slice of the 3D feature volume, enhancing the long-range modelling capability of their model.
+OCCFormer [15] introduces a dual-path transformer block to refine the BEV slice of the 3D feature volume, enhancing the long-range modeling capability of their model.
 FB-OCC [19] proposes an additional backward view-transformation module to improve the semantic information in the final 3D feature volume.
 Multi-Scale Occ [18] leverages a multi-scale fusion mechanism to capture global and local detail information in the 3D feature volume.
-While depth-estimation-based approaches have achieved remarkable performance, they have one significant drawback: requiring depth ground truth labels to boost depth-estimation performance, which in turn boosts the model’s overall performance.
+While depth-estimation-based approaches have achieved remarkable performance, they have one significant drawback: requiring depth ground truth labels to boost depth-estimation performance, boosting the model’s overall performance.
 This requirement introduces extra effort during the training process.
 In this paper, we propose a method that eliminates the need for depth estimation by solely relying on multi-view images to construct a 3D feature volume.
-Our approach achieves superior performance compared to depth-estimation based approaches.
-II-C Query-Based BEV Perception
-The BEVFormer series [26, 47] is one of the pioneering query-based approaches to BEV perception.
-By defining a BEV query plane and utilizing intrinsic and extrinsic parameters, the query vector in each grid of the BEV plane can be projected onto multi-view images, enabling the aggregation of features and the generation of the final BEV feature.
-The temporal self-attention module also leverages historical BEV features to enhance the model’s performance further.
-Similarly, the DETR3D [48] predefines a set of 3D query points as reference points.
-The intrinsic and extrinsic parameters are then used to project those points onto multi-view images to aggregate features.
-These works successfully eliminate the need for depth estimation by employing transformer blocks and relying solely on multi-view image features for BEV perception.
-II-D Query-Based 3D Semantic Occupancy Prediction
+Our approach achieves superior performance compared to depth-estimation-based approaches.
+II-B Query-Based 3D Semantic Occupancy Prediction
 Building upon the success of query-based BEV perception algorithms, TPVFormer [13] introduces an extension to the BEV query by encompassing three perpendicular planes.
@@
 In our case, a class value of 0 indicates that the grid is empty.
-Figure 2 exhibits the overall architecture of our method.
+Fig.
+2 exhibits the overall architecture of our method.
 Initially, given a set of surrounding multi-camera images, we first use a 2D backbone network(e.g.ResNet101-DCN) to extract $Nc$ cameras and $L$ levels multi-scale features $X=\left\{\left\{X_{n}^{l}\right\}_{n=1}^{Nc}\in R^{C_{l}\times H_{l}\times W_{l}}\right\}_{l=1}^{L}$.
 For each level, we construct two projection matrices, namely, global projection matrix $VT\_XY^{l}\in R^{(Nc\times H_{l}\times W_{l})\times(X_{l}\times Y_{l})}$ and local projection matrix $VT\_XYZ^{l}\in R^{(Nc\times H_{l}\times W_{l})\times(X_{l}\times Y_{l}\times Z_{l})}$.
-The feature maps at each level are multiplied with these projection matrices, resulting in the 3D feature volume $F_{local}^{l}\in{R}^{C_{l}\times X_{l}\times Y_{l}\times Z_{l}}$ and the Bird’s-eye view (BEV) feature $F_{global}^{l}\in{R}^{C_{l}\times X_{l}\times Y_{l}}$.
+The feature maps at each level are multiplied with these projection matrices, resulting in the 3D feature volume $F_{local}^{l}\in\mathbb{R}^{C_{l}\times X_{l}\times Y_{l}\times Z_{l}}$ and the Bird’s-eye view (BEV) feature $F_{global}^{l}\in\mathbb{R}^{C_{l}\times X_{l}\times Y_{l}}$.
 Subsequently, the global-local attention fusion module merges information from these two features, producing the final 3D volume.
@@
 III-B Global and Local Projection Matrix Generation
-Figure 3: The predefined 3D volume feature sampling process.
-Each voxel grid is initially divided into $N^{3}$ subspaces along the vertical and horizontal directions, and the centre of each subspace serves as a sample point.
-These sample points are then projected onto all the multi-view feature maps to aggregate the corresponding features for the voxel grid.
 In our approach, constructing global and local projection matrices is critical in gathering information for the local 3D feature volume and the global BEV feature.
 This differs from the method used in Occformer[15], which relies on depth estimation and voxel pooling, or the methods employed in TPVFormer[13] and SurroundOcc[14], which involve transformer-based query.
-To begin with, we establish a set of predefined 3D volume spaces denoted by $F_{3D}^{l}\in{R}^{X^{l}\times Y^{l}\times Z^{l}}$ for each level of multi-view feature maps under the ego vehicle’s coordinate system.
+Figure 3: The predefined 3D volume feature sampling process.
+Each voxel grid is initially divided into $N^{3}$ subspaces along the vertical and horizontal directions, and the center of each subspace serves as a sample point.
+These sample points are then projected onto all the multi-view feature maps to aggregate the corresponding features for the voxel grid.
+To begin with, we establish a set of predefined 3D volume spaces denoted by $F_{3D}^{l}\in\mathbb{R}^{X^{l}\times Y^{l}\times Z^{l}}$ for each level of multi-view feature maps under the ego vehicle’s coordinate system.
 The ego vehicle is positioned at the center of these 3D volumes.
@@
 Sample points that fall outside the boundaries of the feature maps or generate negative depth values are filtered out.
-The features hit by each sample point are then aggregated, resulting in a feature vector representing the corresponding voxel grid (see Figure 3).
-The feature sampling process is a static mapping and can be represented by constructing the projection matrices $VT\_XYZ^{l}\in R^{(Nc\times H_{l}\times W_{l})\times(X_{l}\times Y_{l}\times Z_{l})}$ as exhibited in Figure 4.
-Moreover, height information can be compressed further by constructing the projection matrix $VT\_XY^{l}\in R^{(Nc\times H_{l}\times W_{l})\times(X_{l}\times Y_{l})}$ as shown in Figure 5.
+The features hit by each sample point are then aggregated, resulting in a feature vector representing the corresponding voxel grid (see Fig.
+3).
+The feature sampling process is a static mapping and can be represented by constructing the projection matrices $VT\_XYZ^{l}\in R^{(Nc\times H_{l}\times W_{l})\times(X_{l}\times Y_{l}\times Z_{l})}$ as exhibited in Fig.
+4(a).
+Moreover, height information can be compressed further by constructing the projection matrix $VT\_XY^{l}\in R^{(Nc\times H_{l}\times W_{l})\times(X_{l}\times Y_{l})}$ as shown in Fig.
+4(b).
 As a result, the generation of both the local 3D feature volume and global BEV feature can be greatly simplified as a matrix multiplication between the multi-view feature maps and the two projection matrices:
@@
 By applying this technique, we can dramatically decrease the GPU memory usage for our highest 3D volume resolution from 15GB to 200MB.
-Figure 4: The local projection matrix VT_XYZ represents the static mapping relationship for the feature sampling process of the local 3D feature volume.
-Figure 5: The global projection matrix VT_XY represents the static mapping relationship for the feature sampling process of the global Bird’s Eye View (BEV) feature.
+(a) Local projection matrix
+(b) Global projection matrix
+Figure 4: (a) The local projection matrix VT_XYZ represents the static mapping relationship for the feature sampling process of the local 3D feature volume.
+(b) The global projection matrix VT_XY represents the static mapping relationship for the feature sampling process of the global Bird’s Eye View (BEV) feature.
 III-C Global Local Attention Fusion
 To enhance the ability of the final 3D feature volume to capture both global and local details, we introduce the Global-Local Attention Fusion module.
-The detailed structure of the Global-Local Attention Fusion module is depicted in Figure 6.
+The detailed structure of the Global-Local Attention Fusion module is depicted in Fig.
+5.
 Figure 5: The global-local fusion module.
@@
 Inspired by [23], the module begins by applying traditional 2D and 3D convolution operations to enhance the global BEV feature and the local 3D feature volume.
-Building on recent advancements that emphasize the significance of locality and efficiency in transformers [52, 53] and the importance of the BEV plane [40], we incorporate an efficient window attention module from [52] and a bottleneck ASPP module to further refine the global BEV feature, enabling it to capture long-range global semantic information.
+Building on recent advancements that emphasize the significance of locality and efficiency in transformers and the importance of the BEV plane [15], we incorporate an efficient window attention module from [24] and a bottleneck ASPP module to further refine the global BEV feature.
 Simultaneously, the local 3D feature volume is processed through the Feed-Forward Network (FFN) to generate attention weights.
@@
 Additionally, the Expand operation refers to expanding the dimension of $F_{global}^{l}$ to match the dimension of $F_{local}^{l}$.
+Method Backbone Params Resolution IoU mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const.
+veh.
+$\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive.
+surf.
+$\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation MonoScene[25] ResNet101-DCN - $200\times 200\times 16$ 23.96 7.31 4.03 0.35 8.00 8.04 2.90 0.28 1.16 0.67 4.01 4.35 27.72 5.20 15.13 11.29 9.03 14.86 Atlas* [26] - - $200\times 200\times 16$ 28.66 15.00 10.64 5.68 19.66 24.94 8.90 8.84 6.47 3.28 10.42 16.21 34.86 15.46 21.89 20.95 11.21 20.54 BEVFormer*[7] ResNet101-DCN 59M $200\times 200$ 30.50 16.75 14.22 6.58 23.46 28.28 8.66 10.77 6.64 4.05 11.20 17.78 37.28 18.00 22.88 22.17 13.80 22.21 TPVFormer[13] ResNet101-DCN 69M $200\times 200\times 16$ 11.51 11.66 16.14 7.17 22.63 17.13 8.83 11.39 10.46 8.23 9.43 17.02 8.07 13.64 13.85 10.34 4.90 7.37 TPVFormer* ResNet101-DCN 69M $200\times 200\times 16$ 30.86 17.10 15.96 5.31 23.86 27.32 9.79 8.74 7.09 5.20 10.97 19.22 38.87 21.25 24.26 23.15 11.73 20.81 C-CONet*[27] ResNet101 118M $200\times 200\times 16$ 26.10 18.40 18.60 10.00 26.40 27.40 8.60 15.70 13.30 9.70 10.90 20.20 33.00 20.70 21.40 21.80 14.70 21.30 LMSCNet*[28] - - $200\times 200\times 16$ 36.60 14.90 13.10 4.50 14.70 22.10 12.60 4.20 7.20 7.10 12.20 11.50 26.30 14.30 21.10 15.20 18.50 34.20 L-CONet*[27] - $200\times 200\times 16$ 39.40 17.70 19.20 4.00 15.10 26.90 6.20 3.80 6.80 6.00 14.10 13.10 39.70 19.10 24.00 23.90 25.10 35.70 SurroundOcc*[14] ResNet101-DCN 180M $200\times 200\times 16$ 31.49 20.30 20.59 11.68 28.06 30.86 10.70 15.14 14.09 12.06 14.38 22.26 37.29 23.70 24.49 22.77 14.89 21.86 InverseMatrixVT3D* (Base) ResNet101-DCN 67M $200\times 200\times 16$ 31.85 18.88 18.39 12.46 26.30 29.11 11.00 15.74 14.78 11.38 13.31 21.61 36.30 19.97 21.26 20.43 11.49 18.47 InverseMatrixVT3D* (Post-Fix) ResNet101-DCN 67M $200\times 200\times 16$ 26.79 15.81 16.47 10.27 21.28 28.29 8.32 13.29 12.90 8.41 10.96 18.49 32.43 11.79 18.27 15.45 9.87 16.52 InverseMatrixVT3D* (Pre-Fix) ResNet101-DCN 67M $200\times 200\times 16$ 31.30 18.42 18.06 12.72 25.99 28.00 10.15 15.98 14.31 10.61 12.49 20.58 35.61 19.40 21.00 20.30 11.06 18.59
+TABLE I: 3D semantic occupancy prediction results on nuScenes validation set. * means method is trained with dense occupancy labels from [14].
+Method IoU mIoU road sidewalk parking other-ground building car truck bicycle motorcycle other-vehicle vegetation trunk terrain person bicyclist motorcyclist fence pole traf.sign LMSCNet [28] 28.61 6.70 40.68 18.22 4.38 0.00 10.31 18.33 0.00 0.00 0.00 0.00 13.66 0.02 20.54 0.00 0.00 0.00 1.21 0.00 0.00 AICNet [29] 29.59 8.31 43.55 20.55 11.97 0.07 12.94 14.71 4.53 0.00 0.00 0.00 15.37 2.90 28.71 0.00 0.00 0.00 2.52 0.06 0.00 3DSketch [30] 33.30 7.50 41.32 21.63 0.00 0.00 14.81 18.59 0.00 0.00 0.00 0.00 19.09 0.00 26.40 0.00 0.00 0.00 0.73 0.00 0.00 JS3C-Net [31] 38.98 10.31 50.49 23.74 11.94 0.07 15.03 24.65 4.41 0.00 0.00 6.15 18.11 4.33 26.86 0.67 0.27 0.00 3.94 3.77 1.45 MonoScene [25] 36.86 11.08 56.52 26.72 14.27 0.46 14.09 23.26 6.98 0.61 0.45 1.48 17.89 2.81 29.64 1.86 1.20 0.00 5.84 4.14 2.25 TPVFormer [13] 35.61 11.36 56.50 25.87 20.60 0.85 13.88 23.81 8.08 0.36 0.05 4.35 16.92 2.26 30.38 0.51 0.89 0.00 5.94 3.14 1.52 InverseMatrixVT3D (Base) 36.22 11.81 52.99 25.84 20.04 0.09 13.17 24.08 10.25 1.85 2.65 6.80 16.98 3.09 27.77 4.01 3.13 0.00 4.94 4.05 2.67
+TABLE II: 3D semantic scene completion performance on SemanticKITTI validation set.
 IV Experimental Results
 IV-A Implementation Details
-The InverseMatrixVT3D uses ResNet101-DCN [54, 55] as a 2D backbone with a checkpoint from FCOS3D [56] to extract image features.
-The features of stages 1,2,3 of the backbone are fed to FPN [57], resulting in 3-level multi-scale image features.
+The InverseMatrixVT3D uses ResNet101-DCN [32] as a 2D backbone with a checkpoint from FCOS3D [33] to extract image features.
+The features of stage 1,2,3 of the backbone are fed to FPN [34], resulting in 3-level multi-scale image features.
 The network architecture comprises four levels ($L=4$), with no skip connection applied to the highest level.
@@
 The learning rate is decayed using a multi-step scheduler.
-For the 3D semantic occupancy prediction task, following the approach presented in [39, 37], we incorporate focal loss [58], Lovasz-softmax loss [59] and scene-class affinity loss [37] to handle the significant sparsity of free space in the ground truth dataset.
 Regarding data augmentation, random resize, rotation and flip operations are applied in the image space, following established practices for BEV-based 3D object detection [5, 6, 35, 7] and the compared methods [13, 14, 15, 25].
-Our model includes two resolution settings.
-The predicted occupancy for evaluation is specified at a resolution of $256\times 256\times 32$ and $200\times 200\times 16$, respectively.
-We further fixed the projection matrices under two resolution settings during testing to make the comparison.
-The model is trained using eight A10G GPUs with 24GB of memory, and it has been trained for 24 epochs on 3D semantic occupancy prediction tasks.
-IV-B Dataset
+The predicted occupancy has a resolution of $200\times 200\times 16$ for full-scale evaluation.
+The model is trained to utilize eight A10 GPUs with 24GB of memory, and it has been trained for 2 days.
+IV-B Loss Function
+To train the model, we use focal loss [36], Lovasz-softmax loss [37] and scene-class affinity loss [25] to handle the significant sparsity of free space in the ground truth dataset.
+The final loss is composed of:
+$Loss=L_{focal}+L_{lovasz}+L_{scal}^{geo}+L_{scal}^{sem}$ (5)
+IV-C Dataset
 The nuScenes dataset [1], a vast autonomous driving dataset, serves as the data source for our experiments.
@@
 We set the range for occupancy prediction as [-50, 50] meters for the X and Y axes and [-5, 3] meters for the Z axis.
-The input images have a $1600\times 900$ pixels resolution, while the final output occupancy is represented with a resolution of $256\times 256\times 32$ for the base version and $200\times 200\times 16$ for the small base version.
+The input images have a $1600\times 900$ pixels resolution, while the final output occupancy is represented with a resolution of $200\times 200\times 16$ for the base version.
 We have conducted experiments on 3D semantic occupancy prediction tasks to provide quantitative results.
@@
 Additionally, we provide qualitative visualizations of the results for the 3D semantic occupancy prediction task.
-IV-C Performance Evaluate Metrics
+To enhance the demonstration of the effectiveness of our approach, we conducted a monocular semantic scene completion experiment on the SemanticKITTI dataset [2] employing the left RGB camera.
+SemanticKITTI contains annotated outdoor LiDAR scans with 21 semantic labels.
+The input image resolution is $1241\times 376$, and the ground truth is voxelized into a grid of dimensions 256x256x32 with a voxel size of 0.2m.
+The evaluation of our model is performed on the validation set.
+IV-D Performance Evaluate Metrics
 To assess the performance of various state-of-the-art (SOTA) algorithms and compare them with our approach in the 3D semantic occupancy prediction task, we utilize the intersection over union (IoU) to evaluate each semantic class.
@@
 where $TP$, $FP$, and $FN$ represent the counts of true positives, false positives, and false negatives in our predictions, respectively, while $Cls$ denotes the total class number.
-IV-D Model Performance Analysis
-We perform the task of multi-camera 3D semantic occupancy prediction on the nuScenes dataset.
-To evaluate the performance of our proposed InverseMatrixVT3D, we compare it with SOTA algorithms and present the results in Table I.
-All methods are trained using dense 3D occupancy ground truth labels obtained from [39].
-Our model has two different settings for two resolutions: the InverseMatrixVT3D(Base) and InverseMatrixVT3D(Small Base), which includes the projection matrices generation process, and the InverseMatrixVT3D(Fix) and InverseMatrixVT3D(Small Fix), which fixes the projection matrices.
-Our model displays remarkable performance in both settings in predicting crucial dynamic objects on the road, including buses, cars, construction vehicles, and trucks.
-Additionally, our model excels in identifying diverse background objects, such as drivable surfaces, terrain, and vegetation.
-These results unequivocally highlight our model’s superior 3D world modelling capability compared to other SOTA algorithms.
-Our model aggregates features for the final 3D volume based on sampling at specific sampling locations.
+Refer to captionRefer to caption
+$\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ construction vehicle $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck
+$\bullet$ driveable surface $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation $\bullet$ ego vehicle
+Figure 6: Challenging scenes qualitative analysis.
+Despite challenging lighting and weather conditions, our approach successfully predicts moving objects with high accuracy, even when they are far away from the ego vehicle.
+Better viewed when zoomed in.
+IV-E Model Performance Analysis
+We perform the task of multi-camera 3D semantic occupancy prediction on the nuScenes dataset and the monocular semantic scene completion task on the SemanticKITTI dataset.
+To evaluate the performance of our proposed InverseMatrixVT3D in each task, we compare it with other SOTA algorithms and present the results in Table I and II, respectively.
+In Table I, our model has three different settings: the InverseMatrixVT3D (Base), where the projection matrices generation process is included during training and evaluation, the InverseMatrixVT3D (Post-Fix), which involves fixing the projection matrices after the model finished training for evaluation purposes, and the InverseMatrixVT3D (Pre-Fix), which fixed projection matrices in advance and directly incorporating them into the training process.
+In the context of the 3D semantic occupancy prediction task using the nuScenes dataset, our model demonstrates very competitive performance.
+It outperforms several transformer-query-based methods and ranks second on the benchmark.
+Compared to transformer-query-based methods, our model exhibits the best performance in detecting VRU on roadways, including bicycles, motorcycles, and pedestrians.
+Nevertheless, when dealing with background objects, like vegetation, manmade structures, and terrain, our model is inferior to transformer-query-based approaches.
+This may be attributed to the fact that many background objects are invisible, but due to the powerful generalization ability provided by the transformer, those transformer-query-based approaches can infer the background object’s semantic label and its locations to some extent.
+Another point worth mentioning is that although our model failed to beat SurroundOcc, the size of our model only has a total of 67M trainable parameters, which is substantially smaller than SurroundOcc’s 180M trainable parameters.
+Notably, the model’s performance significantly diminishes when utilizing fixed projection matrices after training due to noticeable projection errors.
+The projection error can be alleviated by employing fixed projection matrices in advance and directly incorporating them into the training process.
+In the realm of the monocular semantic scene completion task on the SemanticKITTI dataset, our model exhibits similar characteristics to those observed in the 3D semantic occupancy prediction task and has delivered very competitive performance in comparison with other SOTA algorithms.
+These results highlight our model’s superior 3D world modelling capability.
+In general, our model aggregates features for the final 3D volume based on sampling at specific sampling locations.
 Compared to the transformer-based approach, whose sampling locations can be adjusted based on the query vector, our approach’s sampling locations are fixed.
-Thus, when dealing with some small objects such as traffic cones and bicycles, our method aggregates fewer features from the image and is inferior to SurroundOcc in predicting those classes.
-However, our sampling density is significantly larger than the transformer-based approach, enabling our method to make good predictions on relatively large objects in the scene.
-In summary, our method sacrifices sampling flexibility to improve the sampling density.
-This underlying principle adjustment boosts model performance significantly in predicting relatively large objects in the scene.
-Method Resolution 3DOcc mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const.
-veh.
-$\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive.
-surf.
-$\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation MonoScene [37] $200\times 200\times 16$ 7.31 4.03 0.35 8.00 8.04 2.90 0.28 1.16 0.67 4.01 4.35 27.72 5.20 15.13 11.29 9.03 14.86 Atlas [60] $200\times 200\times 16$ 15.00 10.64 5.68 19.66 24.94 8.90 8.84 6.47 3.28 10.42 16.21 34.86 15.46 21.89 20.95 11.21 20.54 BEVFormer [26] $200\times 200$ 16.75 14.22 6.58 23.46 28.28 8.66 10.77 6.64 4.05 11.20 17.78 37.28 18.00 22.88 22.17 13.80 22.21 TPVFormer [38] $200\times 200\times 16$ 17.10 15.96 5.31 23.86 27.32 9.79 8.74 7.09 5.20 10.97 19.22 38.87 21.25 24.26 23.15 11.73 20.81 SurroundOcc [39] $200\times 200\times 16$ 20.30 20.59 11.68 28.06 30.86 10.70 15.14 14.09 12.06 14.38 22.26 37.29 23.70 24.49 22.77 14.89 21.86 InverseMatrixVT3D (Base) $256\times 256\times 32$ 26.03 26.15 9.76 42.27 41.89 17.83 14.74 22.88 10.30 21.78 36.90 50.82 21.15 24.45 26.76 16.86 32.03 InverseMatrixVT3D (Small Base) $200\times 200\times 16$ 25.95 26.13 15.19 42.65 43.92 11.42 20.82 22.68 13.45 22.33 36.04 47.47 16.26 21.47 25.33 16.37 33.73 InverseMatrixVT3D (Fix) $256\times 256\times 32$ 24.58 23.19 7.86 43.17 43.16 16.80 16.03 18.16 8.81 24.67 35.93 41.60 17.68 21.11 21.54 22.06 31.45 InverseMatrixVT3D (Small Fix) $200\times 200\times 16$ 22.68 20.91 12.06 38.52 35.53 10.53 21.49 20.12 12.36 20.49 30.31 36.23 14.60 19.25 22.78 15.89 31.76
-TABLE I: 3D semantic occupancy prediction results on nuScenes validation set. All methods are trained with dense occupancy labels from [39].
-IV-E Challenging Scenes Qualitative Analysis
-Figure 6a presents a scene where multiple objects of different semantic classes overlap and crowd.
-Despite this complexity, our method accurately predicts each object’s semantic class and 3D spatial location.
-Furthermore, the edges between objects of different classes are delineated.
-Figure 6b exhibits the scene where a construction vehicle is severely obscured by a large truck while multiple traffic cones and barriers crowd and overlap.
-Even under such a complex scenario, our model can still infer the general occupancy status of the construction vehicle and generate a meticulous 3D occupancy prediction with distinct boundaries.
-We further demonstrate our model’s powerful 3D modelling capability by presenting the prediction results in challenging rainy and night-time scenes, as shown in Figure 8.
-In each subfigure, the prediction result generated by our model is displayed on the bottom left, while the ground truth is shown on the bottom right.
-Regarding the rainy scenario depicted in Figure 7a, we observed an inconsistency in the ground truth labels.
+Nevertheless, our method boasts a significantly higher sampling density than the transformer-based approach, allowing for dense feature aggregation.
+In essence, our model prioritizes sampling density over sampling flexibility, which leads to competitive performance and a much smaller model size.
+IV-F Challenging Scenes Qualitative Analysis
+We demonstrate our model’s powerful 3D modelling capability and exceptional VRU detection performance by presenting the prediction results in challenging rainy and night-time scenes along with SurroundOcc prediction results and ground truth, as shown in Fig.
+6.
+Regarding the rainy scenario depicted in Fig.
+6 upper, we observed an inconsistency in the ground truth labels.
 This discrepancy is attributed to the lidar’s poor performance during rainy weather due to the reflection effect.
@@
 It not only fills in the missing information from the ground truth but also successfully detects a pedestrian crossing the road (indicated by the red circle), even though the person is located at a significant distance from the ego vehicle.
-Moreover, in the night-time scene illustrated in Figure 7b, our model effectively predicts the presence of a motorcycle (marked by the red circle) in the distance ahead despite the low ambient light conditions.
-We also present scenes showcasing crowded multiple objects and rainy day crossroads in Figure 9.
-The left column displays the multi-camera images, the middle column shows the prediction results of our model, and the right column exhibits the ground truth labels.
-In the scenarios with crowded multiple objects represented in the first and second rows of Figure 9, our model accurately segments each object.
-Furthermore, in the rainy city intersection scenario depicted in the last row of Figure 9, our model successfully predicts the road structure even in the presence of missing information in the ground truth labels.
-a Highly overlapping scene
-b Severely obscured scene
-$\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ construction vehicle $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck
-$\bullet$ driveable surface $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation $\bullet$ ego vehicle
-Figure 7: Challenging scenes qualitative analysis.
-Despite the presence of multiple objects that overlap or severely obscure each other, our method is capable of accurately predicting detailed occupancy and delineating their edges.
-Better viewed when zoomed in.
-a Rainy scene
-b Night-time scene
-$\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ construction vehicle $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck
-$\bullet$ driveable surface $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation $\bullet$ ego vehicle
-Figure 8: Rainy and night-time scene.
-Despite challenging lighting conditions, our approach successfully predicts moving objects with high accuracy, even when they are far away from the ego vehicle.
-Better viewed when zoomed in.
-a Multi-Camera Images
-b Prediction Result
-c Ground-Truth
-$\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ construction vehicle $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck
-$\bullet$ driveable surface $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation $\bullet$ ego vehicle
-Figure 9: Challenge scenes. Despite multiple different objects being crowded or missing information in the ground truth, our model can still accurately segment each object.
-IV-F Model Efficiency
-Method Latency (s) ($\downarrow$) Memory (GB) ($\downarrow$) NeWCRFs [61] 1.07 14.5 MonoScene [37] 0.87 20.3 Adabins [62] 0.75 15.5 SurroundDepth [18] 0.73 12.4 SurroundOcc [39] 0.34 5.9 TPVFormer [38] 0.32 5.1 BEVFormer [26] 0.31 4.5 InverseMatrixVT3D (Base) 0.68 10.2 InverseMatrixVT3D (Small Base) 0.5 7.8 InverseMatrixVT3D (Fix) 0.45 4.88 InverseMatrixVT3D (Small Fix) 0.34 4.82
-TABLE II: Model efficiency comparison of different methods.
+Moreover, in the night-time scene illustrated in Fig.
+6 bottom, our model effectively predicts the presence of a motorcycle (marked by the red circle) in the distance ahead despite the low ambient light conditions.
+IV-G Model Efficiency
+Method Latency (s) ($\downarrow$) Memory (GB) ($\downarrow$) NeWCRFs [38] 1.07 14.5 MonoScene [25] 0.87 20.3 Adabins [39] 0.75 15.5 SurroundDepth [3] 0.73 12.4 SurroundOcc [14] 0.34 5.9 TPVFormer [13] 0.32 5.1 BEVFormer [7] 0.31 4.5 InverseMatrixVT3D (Base) 0.5 5.2 InverseMatrixVT3D (Fix) 0.32 4.82
+TABLE III: Model efficiency comparison of different methods.
 The experiments are performed on a single RTX 3090 using six multi-camera images.
 For input image resolution, all methods adopt $1600\times 900$.
-$\downarrow$:the lower the better.
-Table II presents a comparison of the inference time and inference memory among different methods.
+$\downarrow$:the lower, the better.
+Table III compares the inference time and inference memory among different methods.
 The experiments are conducted on a single RTX 3090 using six multi-camera images.
 All methods adopt an image resolution of $1600\times 900$.
-The two base version models, which include the generation of projection matrices, have runtimes of 0.68 and 0.5 seconds for a single data sample.
-However, by fixing the projection matrices and removing the generation procedure, the fixed version model achieves exceptional real-time performance due to the utilization of projection matrices that facilitate 3D volume generation.
-IV-G Ablation Study
-IV-G1 Global Local Attention Fusion
-We perform an ablation study on the global-local attention fusion module, and the experiment results are presented in Table III.
-When we remove the fusion module and directly use the local 3D feature volume as the final 3D volume, we observe around 1.9% degradation in the model’s performance.
-This degradation could be attributed to the inability of the final 3D volume to capture long-range global semantic information.
-Method 3DOcc mIoU w fusion module 23.19 w/o fusion module 21.32
-TABLE III: Ablation study for global-local attention fusion module
-IV-G2 Multi-Scale Supervision
-We conducted an ablation study on the multi-scale supervision mechanism, and the experiment results are presented in Table IV.
-The upsample and skip connection structure remains in the model where the multi-scale supervision mechanism is removed, but we only utilize the highest-level 3D feature volume for supervision.
-The experimental results demonstrate that the coarse-to-fine supervision schema can enhance performance by approximately 2.2%.
-Method 3DOcc mIoU w Multi-Scale Supervision 23.19 w/o Multi-Scale Supervision 21.07
-TABLE IV: Ablation study for Multi-Scale Supervision
-IV-G3 Dividing Schemas For Each Level
-We performed an ablation study on the division schemas for each level, and the experiment results are presented in Table V.
+Our base model, which includes the projection matrices generation, runs 0.5 seconds for a single data sample, but if we fixed the projection matrices, remove the projection matrices generation procedure.
+Due to the utilization of projection matrices that facilitate the 3D volume generation, our fixed version model achieves exceptional real-time performance.
+IV-H Ablation Study
+IV-H1 Global Local Attention Fusion
+We perform an ablation study on the global-local attention fusion module, and the experiment results are presented in Table IV.
+The experimental results confirm that the BEV feature and its associated refinement procedures are crucial in improving model performance.
+Without the BEV feature, the final 3D volume is unable to capture long-range global semantic information, ultimately leading to performance degradation.
+[…]
+TABLE IV: Ablation study for global-local attention fusion module. Bott.ASPP: bottleneck aspp, Eff.Win.Atten: efficient window attention. $\uparrow$:the higher, the better.
+IV-H2 Multi-Scale Mechanism
+We conducted an ablation study on the multi-scale mechanism, and the experiment results are presented in Table V.
+The experimental results demonstrate that the multi-level supervision and coarse-to-fine refinement structure play a vital role during training.
+Without these two structures, our model experiences a performance degradation of at least 5%.
+[…]
+TABLE V: Ablation study on multi-scale mechanism. Multi.Stru: multi-scale corse-to-fine refinement structure, Multi.Sup: multi-level supervision mechanism. $\uparrow$:the higher, the better.
+IV-H3 Dividing Schemas For Each Level
+We performed an ablation study on the division schemas for each level, and the experiment results are presented in Table VI.
 We used different N settings for levels 1, 2, and 3 to generate sample points, and the experimental results aligned with our expectations.
 Specifically, when the divide setting N is larger, more sample points can be generated, leading to dense sampling and rich feature aggregation.
-We also observed a trend where a larger divide setting at higher levels has a more significant impact on the final model performance, indicating the increased importance of higher-level feature aggregation.
-N setting of level 1,2,3 3DOcc mIoU $[3,4,5]$ 23.19 $[2,3,4]$ 20.01 $[1,2,3]$ 17.24 $[1,4,6]$ 19.33
-TABLE V: Ablation study of dividing schemas for each level
+We also observed a trend where a larger divide setting at higher levels significantly impacts the final model performance, indicating the increased importance of higher-level feature aggregation.
+[…]
+3 4 5 14.44%
+2 3 4 8.45%
+1 2 3 5.66%
+1 4 6 7.73%
+TABLE VI: Ablation study of dividing schemas for each level. $\uparrow$:the higher, the better.
 V Conclusion
-In this paper, we propose InverseMatrixVT3D, a method for vision-centric 3D semantic occupancy prediction.
+In this paper, we propose InverseMatrixVT3D, a vision-centric 3D semantic occupancy prediction method.
 Our approach leverages predefined sample points for each scale of 3D volumes and constructs projection matrices to represent the fixed sampling process.
 Through matrix multiplication between multi-view feature maps and projection matrices in a multi-scale fashion, we generate local 3D feature volumes and global BEV features.
-These features are further merged using our proposed global-local fusion module, resulting in the final 3D volume at each level.
+These features are merged using our proposed global-local fusion module, resulting in the final 3D volume at each level.
 Lastly, the 3D volumes at each level are upsampled and fused using a 3D deconvolution layer.
 Unlike other SOTA algorithms, our approach does not require depth estimation or transformer-based query, making the 3D volume generation process simple and efficient.
-Experimental results on nuScenes datasets demonstrate the superiority of our method in benchmark comparisons.
+Extensive experiments conducted on the nuScenes and SemanticKITTI datasets demonstrate that our method excels in its simplicity and effectiveness and achieves the best performance in detecting VRU for autonomous driving and road safety.